The x component of vector B is -25.0 m and the y component is +40.0 m.

(a) What is the magnitude of vector B?

(b) What is the angle between the vector B and the positive direction of x?
Answer #1

tyu B = xity -asî+ 409 ܡܶܛܽܠ B = Y-40m 547 90° x=-25m #x a) magnitude, B = ( x² + y² -|(-25)+(40)2 B = 47.2m 25 Dinection, Ta